OpenStudy (amber15):

A rectangular table top measures 5/7 of a meter long by 2/7 of a meter wide . what is it's area? my answer is 10/49

OpenStudy (amber15):

wonder if its correct

OpenStudy (ddcamp):

\[\frac{ 5 }{ 7 } * \frac{ 2 }{ 7 } = \frac{ 2*5 }{ 7*7 } = \frac{ 10 }{ 49 }\] Yup, you're right. Just remember that the units are meters².

OpenStudy (amber15):

okay thank you